# Physical insights from the aspect ratio dependence of turbulence in negative triangularity plasmas

### v1

@@ III Aspect ratio scan
Figure 5 displays the results. At conventional aspect ratio, flipping the triangularity does not change the stiffness but rather increases the critical gradient, i.e. the gradient at which modes become unstable. This results in reduced turbulent transport and better confinement properties. It also means that, at fixed heat flux, one can produce a plasma with increased gradients. The observation that, in a standard A tokamak, NT primarily affects the critical gradient and not the stiffness is in agreement with many other works [22, 10]. In contrast, at small A we can see the opposite. Indeed, flipping the triangularity does not alter the critical gradient much but does change the stiffness. In the pITG regime NT reduces the stiffness, while in the pTEM regime NT increases it. All these observations indicate that turbulence responds to a change in shape very differently in conventional tokamaks and STs.
IV Physical picture of ITG at negative triangularity
As shown in section III, the aspect ratio dependence of transport in NT and PT plasmas is complicated. To better understand it, we started by studying the scenario that we expected to be the simplest: the large A pITG case. Indeed, in the limit of $A\gg$1 ITG modes depend only on two main geometry-related quantities: Finite Larmor Radius (FLR) effects and magnetic drifts. We then applied this picture to the conventional and tight A cases, where many other factors arising from geometry can affect turbulence.
IV.1 Impact of magnetic drifts and FLR effects on ITG
In this subsection, we will employ [23] and [24] to show how magnetic drifts and FLR effects contribute to the strength of ITG turbulence. We will start from the standard physical picture of the toroidal ITG instability [23].
Toroidal ITG modes are driven by the magnetic field line curvature and gradients in the magnetic field strength. They can be unstable where the curvature is “bad”, meaning the magnetic field curvature (or equivalently gradient) has a component parallel to the temperature gradient.
We can estimate the characteristic value of this velocity as
$\textbf{v}_{D}=\frac{v_{th}^{2}}{B\Omega}\textbf{b}\wedge\bm{\nabla}B,$ (7)
$v_{E\wedge B}=\frac{\textbf{E}\wedge\textbf{B}}{B^{2}}.$ (8)
One can derive a dispersion relation for the ITG modes, which corresponds to the qualitative physical picture [23]. In gyrofluid theory, it is given by.
$\omega_{D}=\frac{(\tau+b)\eta}{1+3(\tau+b)^{2}}\omega_{*}.$ (12)
This is the optimal condition for the onset of an unstable toroidal ITG. If we set $\tau=1$ and $b=0$ and we recast the formula in terms of velocities, the condition becomes
$\textbf{v}_{D}=\frac{\eta}{4}\textbf{v}_{*}$ (13)
This equation shows that to have the maximum possible growth rate, the magnetic drift and the diamagnetic drift velocities have to be of the same order. Since this is a very simplified model, one should not focus on the actual value of the constant that relates the two velocities. Rather, we should focus on the fact that if the magnetic drift velocity of the ions $\textbf{v}_{D}$ is too high or too low with respect to the diamagnetic velocity $\textbf{v}_{*}$, the growth rate will be small. In simple terms, if the depletion/accumulation zones are traveling too fast or too slow with respect to the ion drift wave, a mismatch between the two develops, making the mode less unstable. The need to satisfy this ”resonance” is key in the explanation of NT beneficial effect on ITGs.
Apart from magnetic drifts, the only other way that the flux surface shape enters into the GK model in the $A\gg 1$ limit is through the FLR effects. Specifically, in the local GK equations, the radial dependence of equilibrium quantities is dropped by the application of a gyroaverage, i.e. an integral average along the fast gyromotion of particles. It can be shown [25] that the gyroaverage of a scalar field $\mathcal{F}$ reduces to a multiplication of the field with the zeroth order Bessel function $J_{0}$, i.e.:
$<\mathcal{F}(\textbf{X})>=\sum_{k_{y},k_{x}}\mathcal{F}(k_{x},k_{y})e^{i\textbf{k}_{\perp}\cdot\textbf{X}}J_{0}(\rho_{\sigma}k_{\perp})$ (14)
IV.2 Impact of plasma shape on magnetic drifts and FLR effects
(a) $\sigma=0.08$
(b) $\sigma=0.3$
(c) $\sigma=1.0$
(d) $\sigma=1.5$
$\textbf{k}_{\perp}=k_{x}\bm{\nabla}x+k_{y}\bm{\nabla}y.$ (15)
Therefore, using the quantities defined in appendix B, we can write $k_{\perp}^{2}$ as
$k_{\perp}^{2}=g^{xx}k_{x}^{2}+2g^{xy}k_{x}k_{y}+g^{yy}k_{y}$ (16)
$\begin{split}\omega_{D}&=\left[\frac{1}{B\Omega}\left(\frac{v_{\perp}^{2}}{2}+v_{\parallel}^{2}\right)\textbf{b}\wedge\bm{\nabla}B\right]\cdot\left[k_{x}\bm{\nabla}x+k_{y}\bm{\nabla}y\right]\\ &=\Lambda\left(k_{x}\underbrace{\left(-\partial_{y}B-\frac{\gamma_{2}}{\gamma_{1}}\partial_{z}B\right)}_{v_{Dx}}+k_{y}\underbrace{\left(\partial_{x}B-\frac{\gamma_{3}}{\gamma_{1}}\partial_{z}B\right)}_{v_{Dy}}\right)\end{split}$ (17)
where the scalar $\Lambda$ comprises all constant terms, $v_{Dx}$ is the radial component of the magnetic drift velocity and $v_{Dy}$ the binormal one. The derivation of this equation is reported in appendix C. Therefore, the geometry of the plasma shape determines the drift frequency through the derivatives of the magnetic field and the elements of the metric tensor. Note that, with the GENE sign convention, a negative value of $\omega_{D}$ corresponds to the bad curvature region and positive values to the good curvature region.
By using Miller local equilibrium geometry specification [13], it is straightforward to calculate $g^{yy}$ and $v_{Dy}$ for any value of triangularity. In figure 7(a) we take the equilibrium of the pITG-2 scenario and progressively change the triangularity. Figure 7(a) shows the flux surface shapes, figures 7(b) and 7(c) show $v_{Dy}$ and $g^{yy}$ as functions of the poloidal angle for different values of triangularity.
We see that, as the triangularity is decreased, the FLR effects become stronger in the vicinity of the outboard midplane, while the magnetic drifts become faster, i.e. more negative. In light of the observations made in the previous section, it is straightforward to say that one mechanism contributing to the stabilization of ITG in NT geometry comes from stronger FLR effects at the outboard midplane, where ITGs are stronger. However, the impact of magnetic drift is not clear. As mentioned before, negative values of $v_{Dy}$ correspond to bad curvature. However, more negative values do not necessarily mean stronger instability. Indeed, as we showed in the previous section, ITG is destabilized by matching the magnetic drift velocity and the diamagnetic drift velocity. Therefore, more negative values of $v_{Dy}$ can imply weaker ITG if $v_{Dy}$ is already larger than the condition set by (10). In the next section we will test this hypothesis with linear and nonlinear simulations.
IV.3 Numerical results
IV.3.1 Linear results - Large $A$
We started with linear GK simulations. The main goal was to verify that the simple gyrofluid model presented in subsections IV.1 and IV.2 holds.
$v_{Dy}(z)=\sigma v_{Dy}^{NT}(z)+v_{Dy0},$ (18)
IV.3.2 Nonlinear results - Large A
Nonlinear GK simulations are much more computationally expensive than linear simulations, which prevents an analogous scan to figure 9. Instead, we artificially swap all the geometric coefficients responsible for FLR effects and magnetic drifts between NT and PT. Indeed, from equations (16) and (17), we can easily identify all the coefficients that one needs to exchange between the two geometries. Clearly, this procedure creates unrealistic equilibria, but it is perfect to isolate the contributions coming from FLR effects and magnetic drifts. Apart from these two effects, everything else is left untouched in the GK equations. A similar exercise was carried out by Merlo and Jenko in [10], but only for linear cases.
Encouraged by these results, we tried to apply our physical picture to the conventional and tight A pITG cases and to other large A pITG cases where other geometric parameters were changed apart from triangularity. In the following subsection, we show the results for the conventional and tight A pITG cases, where we applied the very same methodology showed before. In appendix D we show the results for scans in other geometric parameters (i.e. elongation and magnetic shear).
IV.3.3 Nonlinear results - Conventional and small A
Figure 13: The nonlinear heat fluxes in the (a) pITG-2 and (b) pITG-1 cases with a tight aspect ratio $A=1.7$.
V Physical picture of TEM at negative triangularity
TEMs are driven unstable by a resonance in phase space between the diamagnetic drift velocity of the underlying drift wave and the toroidal precessional drift of trapped electrons . Based on this fundamental drive, one can construct a similar physical picture as we did for ITGs. This has already been done in [11]. It can be shown that the toroidal precessional drift is closely related to the bounce average of the magnetic drift velocity of trapped particles [26].
Figure 14 shows the results. The fifth and sixth color bars show the NT and PT self-consistent cases, i.e. all the physical factors are computed self-consistently with the respective geometry. The fourth column shows the case where only magnetic drifts have been swapped between NT and PT. We can see that the heat fluxes rise, but the increase is mild. If we move further to the left, the next bar corresponds to the case where the FLR effects have been swapped along with magnetic drifts. The heat fluxes are increased further, but the total heat flux is still much lower than in the PT case. The same trend can be observed if we do the same exercise starting from the PT scenario, as shown by the seventh bar. Therefore, we can conclude that, in TEM-dominated regimes, faster magnetic drifts and stronger FLR effects are only some of the factors contributing to the reduction of turbulence by NT. To find the missing pieces of the puzzle, we swapped other geometric factors that could play a role in the stabilization of TEM turbulence. The results correspond to the first and second bar of figure 14. First, we swapped the geometric parameters responsible for the parallel dynamics, i.e. the parallel advection term and the trapping term [27]. The parallel dynamics can play an important role in TEM turbulence because they are related to the bouncing points of the trapped particles. Swapping the parallel dynamics (in addition to the magnetic drifts and FLR effects) is showed by the second column. We observe that this increases the total heat flux, but it is still lower than in the PT case. Therefore, the parallel dynamics are part of the story, but not the most important factor for the stabilization of TEM by NT. We also tried to swap the geometric parameters that enter the computation of the collisional operator (when expressed in parallel velocity $v_{\parallel}$ and magnetic moment $\mu$ coordinates), namely the magnetic field strength and the Jacobian. This effect is important because it regulates the diffusion of particles across the trapping-passing boundary, influencing the fraction of trapped particles and thus TEM turbulence. Accordingly, if we look at the first column, where we have also swapped the collisional diffusion between NT and PT, we can see that it is possible to recover the same heat flux of the self-consistent PT case. Moreover, the change in the heat flux is larger than that caused by all the previous geometrical effects.
For the pTEM case, we can conclude that the physics underlying the confinement benefits of NT is not due to a single physical effect but instead a synergistic combination. This exercise is similar to previous work [11], but the results are slightly different, potentially due to the different equilibria considered. Indeed, [11] was able to recover the heat flux of the opposite triangularity scenario just by swapping magnetic drifts and FLR effects. Here, they explain only part of the picture.
VI SMART
@@
SMART is a novel, compact, spherical tokamak ($1.4\leq A\leq 3.0$) that has recently begun operations at the University of Seville. SMART has a strong geometric flexibility that allows it to produce plasmas with an elongation of $\kappa\leq 3.0$ and a triangularity between $-0.6\leq\delta\leq 0.6$. Both single-null and double-null configurations can be produced. For more technical details, the interested reader is referred to [14, 15, 16, 17].
@@
Given the results of the previous section, experimental results from SMART will be very interesting. All the previous simulations are case studies and can partially predict how a NT ST will behave in the limit of certain idealized regimes. To more realistically model SMART, the input parameters for GK simulations have been obtained from TRANSP predictive simulations for the PT scenario [18]. In figure 15 we show the magnetic equilibria and in figure 16 we show the kinetic profiles predicted by TRANSP for the PT scenario. This is an ohmic heated scenario and, since intrinsic rotation is hard to predict, rotation and flow shear were not considered in the following gyrokinetic simulations.
@@
As already mentioned, one has to be careful in assessing the performance of a scenario using only linear simulations. Indeed, the physics of nonlinear saturation is missing and we cannot compute the final heat fluxes without it. However, a linear analysis can allow one to determine the dominant types of instability as well as to study the linear stability of the modes. We started with a $k_{y}$ scan (at $k_{x}=0$) to determine the type of turbulence that dominates at the ion scale, i.e. $k_{y}\rho_{s}=[0.05-2.0]$. To be consistent with the work shown in previous sections, we did not consider electron scale turbulence (i.e. $k_{y}\rho_{s}\gg 2.0$). However, Electron Temperature Gradient (ETG) modes can play an important role in spherical tokamaks [28, 29]. This is left as future work. The results of the simulations at the ion scale are displayed in figure 17. Comparing the magnitude of the growth rates, we see that NT has larger values than PT. This does not necessarily imply larger nonlinear heat fluxes, but does mean that the modes grow faster. However, the most interesting observation is the nature of turbulence. For the entire ion scale, the spectrum is dominated by MTMs. Indeed, looking at the electrostatic potential $\phi$ and the parallel vector potential $A_{\parallel}$ in figure 18, we can see that these modes have tearing parity.
It is well-known that STs can be dominated by MTM turbulence [30]. MTM turbulence was not significant in the previous sections, as $\beta$ and the ratio of electron to ion temperature gradient were smaller. Therefore, to better characterize MTM turbulence in STs and how it behaves depending on triangularity, we performed parameter scans for the $k_{y}\rho_{i}=0.4$ mode.
@@
Figure 20: The (a) growth rate and (b) real frequency of the $k_{y}\rho_{s}=0.4$ mode as a function of the collisionality $\nu_{C}$ in PT (red) and NT (blue) SMART equilibria.
Figure 21: The (a) growth rate and (b) real frequency of the $k_{y}\rho_{s}=0.4$ mode as a function of the electron temperature gradient $R_{LTe}$ in PT (red) and NT (blue) SMART equilibria.
@@
Next, we carried out a scan in collisionality for the $k_{y}\rho_{i}=0.4$ mode, which is shown in figure 20. Again in this parameter scan we see a mode transition. At exactly zero collisionality, as predicted by previous theoretical work [31], we see that MTM cannot exist and instead the most unstable mode is TEM. For this TEM instability, NT has a lower growth rate than PT. However, as soon as the collisionality is finite, MTM dominates and NT has a faster growth rate. If we then increase collisionality, it reduces the gap between the growth rates of the two geometries. In figure 21, we show a scan of the electron temperature gradient, which is another important drive of the MTM instability. Once again, we can see that PT and NT transition to MTM at somewhat different values, and that once the MTMs become dominant, the modes are more unstable in NT.
@@
This linear GK analysis allowed us to identify the dominant type of turbulence in the PT and NT equilibria expected for SMART. In these equilibria both shapes should be dominated by MTM at the ion scale, and NT should have a stronger linear drive. However, we cannot make strong statements about the real performance of the NT option in SMART. Our linear simulations do not predict the nonlinear saturation of these modes and they did not consider the effect of impurities. However, based on the results obtained in this work, we believe that NT will be most beneficial for SMART in a regime where MTMs are less important and ITGs dominate. For this reason, we performed a triple scan in electron and ion temperature gradients and magnetic shear. The results are shown in figures 23 and 24.

### v2

@@ III Aspect ratio scan
Figure 4 displays the results. At conventional aspect ratio, flipping the triangularity does not change the stiffness but rather increases the critical gradient, i.e. the gradient at which modes become unstable. This results in reduced turbulent transport and better confinement properties. It also means that, at fixed heat flux, one can produce a plasma with increased gradients. The observation that, in a standard $A$ tokamak, NT primarily affects the critical gradient and not the stiffness is in agreement with many other works Merlo et al. (2019); Merlo and Jenko (2023). In contrast, at small $A$ we can see the opposite. Indeed, flipping the triangularity does not alter the critical gradient much but does change the stiffness. In the pITG regime NT reduces the stiffness, while in the pTEM regime NT increases it. All these observations indicate that turbulence responds to a change in shape very differently in conventional tokamaks and STs.
IV SMART
@@
SMART is a novel, compact, spherical tokamak ($1.4\leq A\leq 3.0$) that has recently begun operations at the University of Seville. SMART has a strong geometric flexibility that allows it to produce plasmas with an elongation of $\kappa\leq 3.0$ and a triangularity between $-0.6\leq\delta\leq 0.6$. Both single-null and double-null configurations can be produced. For more technical details, the interested reader is referred to Doyle et al. (2021a, b); Mancini et al. (2023); Agredano-Torres et al. (2021).
@@
Given the results of the previous section, experimental results from SMART will be very interesting. All the previous simulations are case studies and can partially predict how a NT ST will behave in the limit of certain idealized regimes. To more realistically model SMART, the input parameters for GK simulations have been obtained from TRANSP predictive simulations for the PT scenario Cruz-Zabala et al. (2024). In figure 5 we show the magnetic equilibria and in figure 6 we show the kinetic profiles predicted by TRANSP for the PT scenario. This is an ohmic heated scenario and, since intrinsic rotation is hard to predict, rotation and flow shear were not considered in the following gyrokinetic simulations.
@@
As already mentioned, one has to be careful in assessing the performance of a scenario using only linear simulations. Indeed, the physics of nonlinear saturation is missing and we cannot compute the final heat fluxes without it. However, a linear analysis can allow one to determine the dominant types of instability as well as to study the linear stability of the modes. We started with a $k_{y}$ scan (at $k_{x}=0$) to determine the type of turbulence that dominates at the ion scale, i.e. $k_{y}\rho_{s}=[0.05-2.0]$. To be consistent with the work shown in previous sections, we did not consider electron scale turbulence (i.e. $k_{y}\rho_{s}\gg 2.0$). However, Electron Temperature Gradient (ETG) modes can play an important role in spherical tokamaks Guttenfelder and Candy (2011); Mazzucato et al. (2008). This is left as future work. The results of the simulations at the ion scale are displayed in figure 7. Comparing the magnitude of the growth rates, we see that NT has larger values than PT. This does not necessarily imply larger nonlinear heat fluxes, but does mean that the modes grow faster. However, the most interesting observation is the nature of turbulence. For the entire ion scale, the spectrum is dominated by MTMs. Indeed, looking at the electrostatic potential $\phi$ and the parallel vector potential $A_{\parallel}$ in figure 8, we can see that these modes have tearing parity.
It is well-known that STs can be dominated by MTM turbulence Applegate et al. (2007); Dickinson et al. (2012). MTM turbulence was not significant in the previous sections, as $\beta$ and the ratio of electron to ion temperature gradient were smaller. Therefore, to better characterize MTM turbulence in STs and how it behaves depending on triangularity, we performed parameter scans for the $k_{y}\rho_{i}=0.4$ mode.
@@
Figure 10: The (a) growth rate and (b) real frequency of the $k_{y}\rho_{s}=0.4$ mode as a function of the electron-to-ion collisionality $\nu_{ei}$ in PT (red) and NT (blue) SMART equilibria.
Figure 11: The (a) growth rate and (b) real frequency of the $k_{y}\rho_{s}=0.4$ mode as a function of the electron temperature gradient $R/L_{Te}$ in PT (red) and NT (blue) SMART equilibria.
@@
Next, we carried out a scan in collisionality for the $k_{y}\rho_{i}=0.4$ mode, which is shown in figure 10. Again in this parameter scan we see a mode transition. At exactly zero collisionality, as predicted by previous theoretical work Drake and Lee (1977), we see that MTM cannot exist and instead the most unstable mode is TEM. For this TEM instability, NT has a lower growth rate than PT. However, as soon as the collisionality is finite, MTM dominates and NT has a faster growth rate. If we then increase collisionality, it reduces the gap between the growth rates of the two geometries. In figure 11, we show a scan of the electron temperature gradient, which is another important drive of the MTM instability. Once again, we can see that PT and NT transition to MTM at somewhat different values, and that once the MTMs become dominant, the modes are more unstable in NT.
@@
This linear GK analysis allowed us to identify the dominant type of turbulence in the PT and NT equilibria expected for SMART. In these equilibria both shapes should be dominated by MTM at the ion scale, and NT should have a stronger linear drive. However, we cannot make strong statements about the real performance of the NT option in SMART. Our linear simulations do not predict the nonlinear saturation of these modes and they did not consider the effect of impurities. However, based on the results obtained in this work, we believe that NT will be most beneficial for SMART in a regime where MTMs are less important and ITGs dominate. For this reason, we performed a triple scan in electron and ion temperature gradients and magnetic shear. The results are shown in figures 13 and 14.
@@
V Physical picture of ITG at negative triangularity
As shown in sections III and IV, the aspect ratio dependence of transport in NT and PT plasmas is complicated. To better understand it, we started by studying the scenario that we expected to be the simplest: the large $A$ pITG case. Indeed, in the limit of $A\gg$1 ITG modes depend only on two main geometry-related quantities: Finite Larmor Radius (FLR) effects and magnetic drifts. We then applied this picture to the conventional and tight $A$ cases, where many other factors arising from geometry can affect turbulence.
V.1 Impact of magnetic drifts and FLR effects on ITG
In this subsection, we will employ Beer (1994) and Parisi et al. (2020) to show how magnetic drifts and FLR effects contribute to the strength of ITG turbulence.
Starting from the standard physical picture of the toroidal ITG instability Beer (1994), one can derive the relation
$\omega_{D}=\frac{(\tau+b)\eta}{1+3(\tau+b)^{2}}\omega_{*},$ (5)
that determines the condition for which the growth rate $\gamma$ of an ITG mode will be maximized. Here $\omega_{D}$ is the magnetic drift, $\tau=T_{i}/T_{e}$ is the ion to electron temperature ratio, $b=k_{\perp}^{2}\rho_{i}^{2}$, $\rho_{i}$ is the ion Larmor radius, $k_{\perp}$ is the perpendicular wavenumber, $\eta=L_{ni}/L_{Ti}$, where $L_{ni}=-d(\ln{n_{i}})/dx$ is the ion density logarithmic gradient, $L_{Ti}=-d(\ln{T_{i}})/dx$ is the ion temperature logarithmic gradient, $\omega_{*}$ is the diamagnetic drift frequency. All the details regarding this derivation, including the definition of all the mentioned quantities, are reported in appendix B.
If we set $\tau=1$ and $b=0$ and we recast the formula in terms of velocities, the condition becomes
$\textbf{v}_{D}=\frac{\eta}{4}\textbf{v}_{*}.$ (6)
This equation shows that to have the maximum possible growth rate, the magnetic drift and the diamagnetic drift velocities have to be of the same order. Since this is a very simplified model, one should not focus on the actual value of the constant that relates the two velocities. Rather, we should focus on the fact that if the magnetic drift velocity of the ions $\textbf{v}_{D}$ is too high or too low with respect to the diamagnetic velocity $\textbf{v}_{*}$, the growth rate will be small. In simple terms, if the depletion/accumulation zones of ions are traveling too fast or too slow with respect to the ion drift wave, a mismatch between the two develops, making the mode less unstable. The need to satisfy this ”resonance” is key in the explanation of NT beneficial effect on ITGs.
Apart from magnetic drifts, the only other way that the flux surface shape enters into the GK model in the $A\gg 1$ limit is through the FLR effects. Specifically, in the local GK equations, the radial dependence of equilibrium quantities is dropped by the application of a gyroaverage, i.e. an integral average along the fast gyromotion of particles. It can be shown Merz (2008) that the gyroaverage of a scalar field $\mathcal{F}$ reduces to a multiplication of the field with the zeroth order Bessel function $J_{0}$, i.e.:
$<\mathcal{F}(\textbf{X})>=\sum_{k_{y},k_{x}}\mathcal{F}(k_{x},k_{y})e^{i\textbf{k}_{\perp}\cdot\textbf{X}}J_{0}(\rho_{\sigma}k_{\perp}),$ (7)
V.2 Impact of plasma shape on magnetic drifts and FLR effects
$\textbf{k}_{\perp}=k_{x}\bm{\nabla}x+k_{y}\bm{\nabla}y.$ (8)
Therefore, using the quantities defined in appendix C, we can write $k_{\perp}^{2}$ as
$k_{\perp}^{2}=g^{xx}k_{x}^{2}+2g^{xy}k_{x}k_{y}+g^{yy}k_{y}.$ (9)
$\begin{split}\omega_{D}&=\left[\frac{1}{B\Omega}\left(\frac{v_{\perp}^{2}}{2}+v_{\parallel}^{2}\right)\textbf{b}\wedge\bm{\nabla}B\right]\cdot\left[k_{x}\bm{\nabla}x+k_{y}\bm{\nabla}y\right]\\ &=\Lambda\left(k_{x}\underbrace{\left(-\partial_{y}B-\frac{\gamma_{2}}{\gamma_{1}}\partial_{z}B\right)}_{v_{Dx}}+k_{y}\underbrace{\left(\partial_{x}B-\frac{\gamma_{3}}{\gamma_{1}}\partial_{z}B\right)}_{v_{Dy}}\right),\end{split}$ (10)
where the scalar $\Lambda$ comprises all constant terms, $v_{Dx}$ is the radial component of the magnetic drift velocity and $v_{Dy}$ the binormal one. The derivation of this equation is reported in appendix D. Therefore, the geometry of the plasma shape determines the drift frequency through the derivatives of the magnetic field and the elements of the metric tensor. Note that, with the GENE sign convention, a negative value of $\omega_{D}$ corresponds to the bad curvature region and positive values to the good curvature region.
By using Miller local equilibrium geometry specification Miller et al. (1998), it is straightforward to calculate $g^{yy}$ and $v_{Dy}$ for any value of triangularity. In figure 15(a) we take the equilibrium of the pITG-2 scenario and progressively change the triangularity. Figure 15(a) shows the flux surface shapes, figures 15(b) and 15(c) show $v_{Dy}$ and $g^{yy}$ as functions of the poloidal angle for different values of triangularity.
We see that, as the triangularity is decreased, the FLR effects become stronger in the vicinity of the outboard midplane, while the magnetic drifts become faster, i.e. more negative. In light of the observations made in the previous section, it is straightforward to say that one mechanism contributing to the stabilization of ITG in NT geometry comes from stronger FLR effects at the outboard midplane, where ITGs are stronger. However, the impact of magnetic drift is not clear. As mentioned before, negative values of $v_{Dy}$ correspond to bad curvature. However, more negative values do not necessarily mean stronger instability. Indeed, as we showed in the previous section, ITG is destabilized by matching the magnetic drift velocity and the diamagnetic drift velocity. Therefore, more negative values of $v_{Dy}$ can imply weaker ITG if $v_{Dy}$ is already larger than the condition set by equation (6). In the next section we will test this hypothesis with linear and nonlinear simulations.
V.3 Numerical results
V.3.1 Linear results - Large $A$
We started with linear GK simulations. The main goal was to verify that the simple gyrofluid model presented in subsections V.1 and V.2 holds.
$v_{Dy}(z)=\sigma v_{Dy}^{NT}(z)+v_{Dy0},$ (11)
V.3.2 Nonlinear results - Large $A$
Nonlinear GK simulations are much more computationally expensive than linear simulations, which prevents an analogous scan to figure 17. Instead, we artificially swap all the geometric coefficients responsible for FLR effects and magnetic drifts between NT and PT. Indeed, from equations (9) and (10), we can easily identify all the coefficients that one needs to exchange between the two geometries. Clearly, this procedure creates unrealistic equilibria, but it is perfect to isolate the contributions coming from FLR effects and magnetic drifts. Apart from these two effects, everything else is left untouched in the GK equations. A similar exercise was carried out by Merlo and Jenko in Merlo and Jenko (2023), but only for linear cases.
Encouraged by these results, we tried to apply our physical picture to the conventional and tight $A$ pITG cases and to other large $A$ pITG cases where other geometric parameters were changed apart from triangularity. In the following subsection, we show the results for the conventional and tight $A$ pITG cases, where we applied the very same methodology showed before. In appendix E we show the results for scans in other geometric parameters (i.e. elongation and magnetic shear).
V.3.3 Nonlinear results - Conventional and small A
Figure 21: The nonlinear heat fluxes in the (a) pITG-2 and (b) pITG-1 cases with tight aspect ratios $A=2.0$ and $A=1.7$ respectively.
VI Physical picture of TEM at negative triangularity
TEMs are driven unstable by a resonance in phase space between the diamagnetic drift velocity of the underlying drift wave and the toroidal precessional drift of trapped electrons. Based on this fundamental drive, one can construct a similar physical picture as we did for ITGs. This has already been done in Marinoni et al. (2009). It can be shown that the toroidal precessional drift is closely related to the bounce average of the magnetic drift velocity of trapped particles Graves (2013).
Figure 22 shows the results. The fifth and sixth color bars show the NT and PT self-consistent cases, i.e. all the physical factors are computed self-consistently with the respective geometry. The fourth column shows the case where only magnetic drifts have been swapped between NT and PT. We can see that the heat fluxes rise, but the increase is mild. If we move further to the left, the next bar corresponds to the case where the FLR effects have been swapped along with magnetic drifts. The heat fluxes are increased further, but the total heat flux is still much lower than in the PT case. The same trend can be observed if we do the same exercise starting from the PT scenario, as shown by the seventh bar. Therefore, we can conclude that, in TEM-dominated regimes, faster magnetic drifts and stronger FLR effects are only some of the factors contributing to the reduction of turbulence by NT. To find the missing pieces of the puzzle, we swapped other geometric factors that could play a role in the stabilization of TEM turbulence. The results correspond to the first and second bar of figure 22. First, we swapped the geometric parameters responsible for the parallel dynamics, i.e. the parallel advection term and the trapping term Görler (2009). The parallel dynamics can play an important role in TEM turbulence because they are related to the bouncing points of the trapped particles. Swapping the parallel dynamics (in addition to the magnetic drifts and FLR effects) is showed by the second column. We observe that this increases the total heat flux, but it is still lower than in the PT case. Therefore, the parallel dynamics are part of the story, but not the most important factor for the stabilization of TEM by NT. We also tried to swap the geometric parameters that enter the computation of the collisional operator (when expressed in parallel velocity $v_{\parallel}$ and magnetic moment $\mu$ coordinates), namely the magnetic field strength and the Jacobian. This effect is important because it regulates the diffusion of particles across the trapping-passing boundary, influencing the fraction of trapped particles and thus TEM turbulence. Accordingly, if we look at the first column, where we have also swapped the collisional diffusion between NT and PT, we can see that it is possible to recover the same heat flux of the self-consistent PT case. Moreover, the change in the heat flux is larger than that caused by all the previous geometrical effects.
For the pTEM case, we can conclude that the physics underlying the confinement benefits of NT is not due to a single physical effect but instead a synergistic combination. This exercise is similar to previous work Marinoni et al. (2009), but the results are slightly different, potentially due to the different equilibria considered. Indeed, Marinoni et al. (2009) was able to recover the heat flux of the opposite triangularity scenario just by swapping magnetic drifts and FLR effects. Here, they explain only part of the picture.